OpenStudy (squishy):

Solve by elimination. 24x + 2y = 52 6x - 3y = -36 Need Help Please explain

OpenStudy (anonymous):

multiply the bottom equation by -4 so the x's cancel out then add the remaing terms

OpenStudy (anonymous):

24x +2y = 52 6x-3y = -36 ..... multiply this equation with -4 then add both equation we get -24x + 12y = 144. after adding 14y=196 y= 196/14 = 15. so y =15 and x =1.5

OpenStudy (squishy):

ok so where does the four come from?

OpenStudy (anonymous):

just multiply -4 to eliminate 1 variable x here

OpenStudy (squishy):

ook thanks